# A time warping model for seasonal data with application to age estimation from narwhal tusks

### v1

@@ 5 Age determination of a narwhal tusk
Narwhals exhibit a consistent and well-documented annual migration pattern as they move between specific summer and winter habitats, which are situated in fjords and inlets, and offshore areas, respectively (Heide-Jørgensen et al., 2020). The primary prey items consist of halibut, squids, and polar cod (Laidre and Heide-Jørgensen, 2005; Watt and Ferguson, 2014). On their seasonal migration routes, the narwhals travel through the same regional water masses to and from summer- and winter grounds. The strict diet and rigid migration pattern results in distinguishable GLGs, the annual depositions of growth layers in the dentin zone of the tusk, consisting of an opaque and a translucent layer, see Figure 5(b). Continuous fluctuations in the deposition/biomineralization rates show up in the signals and likely occur due to seasonal variations driven by environmental factors, such as temperature or salinity (Hüssy et al., 2021b). In addition, the physiological characteristics of the animals, notably age and health, likely modulate metabolic processes and as a result potentially influence the biomineralization rate (Mosegaard, Svedäng and Taberman, 1988; Hüssy and Mosegaard, 2004). These assumptions are based on otoliths studies, and it seems reasonable to assume that the tusk growth rate exhibit similar characteristics. The recurring deposition of GLGs in conjunction with a variable deposition rate produces a cyclic profile that fits well into our working model (1). Since the hypothesis is that one GLG is deposited annually, counting these cycles will provide an age estimate of the narwhal. An approximate determination of the range of GLGs in a tusk piece can be made through visual observation, see Figure 5(b). However, the visual counts are highly uncertain for some tusk pieces, and automatic robust methods based on more objective criteria and quantitative measures are needed.
@@ 5.1 Data acquisition and preparation
We analyzed one tusk (ID 956) collected from the Inuit Hunt of narwhals in Niaqornat, West Greenland, in 2010. The tusk was sectioned in two halves and one was used for subsequent analysis. The half tusk were divided into several tusk pieces covering all GLGs (see Figure 5(a)). This resulted in $12$ pieces. Each tusk piece was measured along a transect (see Figure 5(b)) using Laser Ablation Inductively Coupled Plasma Mass Spectrometry (LA-ICP-MS) by the Geological Survey of Denmark and Greenland (GEUS), providing the concentrations of 14 isotopes and trace elements (Sr, Ba, Zn, Mg, K, Li, Mn, Pb, Cu, P, Ti, Cr, Co, Rb) (Thomsen et al., 2022; Koch and Günther, 2011). Our analysis centers on Barium-137 normalized by the isotope Calcium-43 in high abundance, since among the set of elements, Barium displayed distinct cyclic patterns speculated to correlate with migration between summer and winter grounds (Heimbrand et al., 2020; Limburg et al., 2011). Consequently two peaks in the signal would correspond to one annual cycle. The combined signal over all pieces displaying the cyclic pattern is visualized for tusk 956 in Figure 6.
Figure 6: Analyzed data set. Calcium-43 adjusted Barium-137 signals obtained from tusk 956. The data set consists of 12 piecewise signals merged together into one common time series in chronological order. The first part of the time series corresponds to the tip of the tusk, deposited when the whale was born, and the last part of the series corresponds to the base nearest the skull, deposited just before death.
@@ 5.2 Dating the observations
For each piece $j$ of the tusk, we fit the model with Algorithm 1 (subpanel A in Figure 7 and Figure 8). We fit each tusk piece individually allowing different parameter values in each, since the values will depend on the angle of the transect line (see red line in Figure 5(b)) across the piece.For example, a steeper line will increase the length of a year, changing the distance scale, which will change all parameters related to the growth process. Likewise, amplitudes are expected to change over the lifetime of the animal, and thus, change from piece to piece. We obtain estimates of the growth time processes $\hat{g}_{j}(x_{ij}),\,i=1,\ldots,n_{j},\,j=1,\ldots,12$, where $n_{j}$ are the number of observations in piece $j$ (Figure 7B). This takes us from the original coordinates $(y_{ij},x_{ij})_{i=1,\ldots,n_{j},j=1,\ldots,12}$ to a new set of coordinates $(y_{ij},t_{ij})$ with $t_{ij}=\hat{g}_{j}(x_{ij})$, where $t_{ij}$ is a increasing sequence proportional to the number of cycles. This shift in coordinates warps the signal and adjust for any non-linearity present in the original time line (Figure 7C).
Figure 7: Time warping of barium signal in narwhal tusk for piece 956-VII. A. Fitted signal (blue curve) and true signal (black curve). B. Estimate $\hat{g}(x)$ of the growth time process $g(x)$. C. Transformation from measured distance $x$ (black curve, same as in panel A) to estimated time $t$ (time warping, magenta curve). Blue dashed lines are the estimated peaks found with equation (6). Magenta dashed lines are the warped peaks. The horizontal axis is displayed both in years (lower axis) and mm (distance on tusk; upper axis).
For each of the 12 pieces, we obtain a growth time process $\hat{g}_{j}(x_{ij})$. The total number of observations is $n_{tot}=\sum_{j=1}^{12}n_{j}$. Re-enumerate the set of indices by concatenating chronologically
$ij\rightarrow m=i+\sum_{l=1}^{j}n_{l},\quad m=1,\ldots,n_{tot}$ (24)
Assuming each piece signal can be glued together, without any gap in the timeline, we can then construct an aggregated growth time process:
$\hat{g}(x_{m})=\hat{g}(x_{ij})=\begin{cases}\hat{g}_{1}(x_{i1})&i\leq n_{1}\\ \hat{g}_{j-1}(x_{n_{j-1},j-1})+\hat{g}_{j}(x_{ij})&n_{j-1}<i\leq n_{j},\,j=2,\ldots,12\end{cases}$ (25)
@@
$\textrm{Year}(m)=2010-\left(\hat{g}(x_{n})-\hat{g}(x_{m})\right)/2\pi,\quad m=1,\ldots,n_{tot}$ (26)
@@ 5.3 Age estimation
The quantity of interest is the age of the narwhal, which is equivalent to the number of cycles easily obtained using formula (5) and $\hat{g}(x_{m})$. A confidence interval is obtained by bootstrapping (Algorithm 4).
This is accomplished by performing parametric bootstrapping on each piece, and combine the bootstrapped signals randomly to produce $M=10^{7}$ realizations of aggregated growth-time processes $\hat{g}(x_{m})$, of which each has its own age estimate.
The age estimate (number of GLGs), including upper and lower confidence bound, are summarized in table 2.
Table 2: Age estimates (total GLGs) with 95% confidence interval. * Based on radiocarbon dating using the bomb pulse (Zoppi et al., 2004) and Carbon-14 measurements from the narwhal tusk (Garde et al., 2024).
@@ 5.4 Model validation
We have overlain the fitted curves with the tusk pieces, to compare the identified peaks with the GLGs of the tusks. In Figure 8 we show two of such examples. The laser line is highlighted in green and red, when inside and outside the dentin area, respectively. The positions of the peaks are marked with vertical red lines, and we can discern how these peaks appear to align with consecutive white lines, which presumably represent one GLG. A more formal model validation was carried out by inspecting the residuals (defined in Section 3.7) visualized in Figure 9. The left plot shows a residual plot (raw residuals against fitted values) for each of the 12 fitted signals and the right column display a QQ-plot, where raw residuals are compared against simulated residuals. The residual plots reveal a skewed pattern for several tusk pieces, likely caused by rapid secondary cycles unique to our case study signals, which are not incorporated in our model (1). See the raw signal in Figure 7 for an illustration of these secondary cycles. For the QQ-plots, most residuals are symmetrically distributed around the diagonal, indicating that the central tendencies are well-captured by the model. However, skewed behavior in the tails, such as in piece Xs, suggests that the model may not fully capture the behavior of extreme values, possibly due to increased variability in regions where the signal has peaks.
@@
Figure 9: Model validation plots for the case-study signals. Residual plots (left) and QQ-plots (right) for all 12 signals of tusk 956.
Estimator uncertainties within each tusk piece is visualized in Figure 10 using violinplots for all parameters. The procedure is described in section 3.7. The simulated signals generally result in consistent estimates for parameters $a,\gamma^{2}$, which centers around the model estimates of the original signals. However, among the $12$ signals used in the case study, Figure 10 also reveals some spurious discoveries. Several estimators display both a sizeable amount of variability and bias. For example, the one-lag autocorrelation $\rho$ is highly variable and for a few signals also show a slight bias. The noise $\sigma$ is drastically underestimated at the early pieces, which exhibit few cycles, and slightly overestimated at later pieces, which contain more cycles. The phase $b$ also displays a high variance, and appears unidentifiable.
@@
Figure 10: Estimator variability of tusk 956. The x-axis represent the different tusk pieces (12 in total). For an estimate $\hat{\theta}$ of some model parameter $\theta$ we compute 100 bootstrap estimates $\tilde{\theta}$ and plot the differences $\Delta\theta:=(\hat{\theta}-\tilde{\theta})/\hat{\theta}$. The difference in phase $\Delta b$ is computed as the shortest distance on the unit circle, not in absolute difference, thus all differences are non-negative.
6 Discussion and research outlook
[…]
@@ 6 Discussion and research outlook
It is important to consider the interplay between frequency and amplitude. For any periodic signal, we need to determine whether changes are primarily driven by variations in amplitude (i.e. magnitude of the signal) or frequency (i.e. rate of the signal). In our approach we construct the varying amplitudes through linear interpolation between peaks while frequency is modeled as a continuous stochastic process. This implies that any instantaneous jumps in the signal are due to changes in frequency, whereas the amplitudes are slowly varying. For the narwhal tusks, this assumption means that rapid changes in the signal occur due to changes in growth rate rather than fluctuations in the uptake of elements in the tusk.
@@
While we believe our presented model to be both flexible and theoretically sound, the practical implementation still requires some choices which should be tailored to the specific signals. Some examples include selection of smoothing parameters in the initialization phase, runtime parameters (number of iterations inside Algorithm 1, particles in SMC, threshold for stopping criteria) and the fraction of neighbors to consider in estimation of the amplitude displacement parameters $r_{k}$’s (see Algorithm 3). These choices do not only assist the method in proper convergence, but are also important for controlling algorithmic runtime. These choices should not be perceived as weaknesses within our model but rather as strategic responses that will serve the algorithm in producing more reliable estimates and guide the optimization procedure.
@@
[Acknowledgments] We would like to thank the Greenlandic hunters for providing the narwhal tusk, Jeppe Møhl for sectioning the tusk, and Tonny Bernt Thomsen and Benjamin Heredia at the Geological Survey of Denmark and Greenland for performing the elemental analyses. The tusk was imported to Denmark from Greenland on CITES import permit no. IM 0822-296/18. Finally, we would like to thank Adam Gorm Hoffman, who was working with modelling and data analysis of a narwhal tusk in the early stage of the project.

### v2

@@ 5 Age determination of a narwhal tusk
First we introduce the case study. The data is presented in Section 5.1. In Section 5.2 we discuss how the model and the data can be used to date the narwhals. In Section 5.3 we provide the estimated age, along with a confidence interval and compare to estimates from manual counts. Section 5.4 presents model validation, checking if the data conform to the model assumptions.
Narwhals exhibit a consistent and well-documented annual migration pattern as they move between specific summer and winter habitats, which are situated in fjords and inlets, and offshore areas, respectively (Heide-Jørgensen et al., 2020). The primary prey items consist of halibut, squids, and polar cod (Laidre and Heide-Jørgensen, 2005; Watt and Ferguson, 2014). On their seasonal migration routes, the narwhals travel through the same regional water masses to and from summer and winter grounds. The strict diet and rigid migration pattern result in annual depositions of growth layers in the dentin zone of the tusk, consisting of an opaque and a translucent layer, see Figure 6(b). Continuous fluctuations in the deposition/biomineralization rates show up in the signals and likely occur due to seasonal variations driven by environmental factors, such as temperature or salinity (Hüssy et al., 2021b). In addition, the physiological characteristics of the animals, notably age and health, likely modulate metabolic processes and as a result potentially influence the biomineralization rate (Mosegaard, Svedäng and Taberman, 1988; Hüssy and Mosegaard, 2004). These assumptions are based on otoliths studies, and it seems reasonable to assume that the tusk growth rate exhibit similar characteristics. The recurring deposition of GLGs in conjunction with a variable deposition rate produces a cyclic profile that fits well into our working model (1). Since the hypothesis is that one GLG is deposited annually, counting these cycles will theoretically provide an age estimate of the narwhal. An approximate determination of the range of GLGs in a tusk piece can be made through visual observation, see Figure 6(b). However, in practice visually reading of the GLGs is both difficult and highly uncertain, and automatic robust methods based on more objective criteria and quantitative measures are needed.
@@ 5.1 Data acquisition and preparation
We analyzed one tusk (ID 956) collected from the Inuit Hunt of narwhals in Niaqornat, West Greenland, in 2010. The tusk was sectioned in two halves and one was used for subsequent analysis. The half tusk was divided into several tusk pieces covering all GLGs (see Figure 6(a)). Few consecutive pieces had a complete overlap of GLGs, thus only one of these were used. After the exclusion, $12$ pieces remained. Each tusk piece was measured along a transect (see Figure 6(b)) using Laser Ablation Inductively Coupled Plasma Mass Spectrometry (LA-ICP-MS) by the Geological Survey of Denmark and Greenland (GEUS), providing the concentrations of 14 isotopes and trace elements (Sr, Ba, Zn, Mg, K, Li, Mn, Pb, Cu, P, Ti, Cr, Co, Rb) (Thomsen et al., 2022; Koch and Günther, 2011). Our analysis centers on Barium-137 normalized by the isotope Calcium-43 in high abundance, since among the set of elements, Barium displayed distinct cyclic patterns speculated to correlate with migration between summer and winter grounds (Heimbrand et al., 2020; Limburg et al., 2011). Consequently two peaks in the signal would correspond to one annual cycle. The combined signal over all pieces is visualized in Figure 7.
Figure 7: Analyzed data set. Calcium-43 adjusted Barium-137 signals obtained from tusk 956. The data set consists of 12 signals merged together into one common time series in chronological order. The first part of the time series corresponds to the tip of the tusk, deposited when the whale was born, and the last part of the series corresponds to the base nearest the skull, deposited just before death.
@@ 5.2 Dating the observations
For each piece $j$ of the tusk, we fit the model with Algorithm 1 (subpanel A in Figure 8 and Figure 9). We fit each tusk piece individually allowing different parameter values in each, since the values will depend on the angle of the transect line (see red line in Figure 6(b)) across the piece. For example, a steeper line will increase the length of a year, changing the distance scale, which will change all parameters related to the growth process. Likewise, amplitudes are expected to change over the lifetime of the animal, and thus, change from piece to piece. We obtain estimates of the growth processes $\hat{g}_{j}(x_{ij}),\,i=1,\ldots,n_{j},\,j=1,\ldots,12$, where $n_{j}$ is the number of observations in piece $j$ (Figure 8B). This takes us from the original coordinates $(y_{ij},x_{ij})_{i=1,\ldots,n_{j},j=1,\ldots,12}$ to a new set of coordinates $(y_{ij},t_{ij})$ with $t_{ij}=\hat{g}_{j}(x_{ij})$, where $t_{ij}$ is an increasing sequence proportional to the number of cycles. This shift in coordinates warps the signal and adjusts for any non-linearity present in the original timeline (Figure 8C).
Figure 8: Time warping of barium signal in narwhal tusk for piece 956-VII. A. Fitted signal (blue curve) and true signal (black curve). B. Estimate $\hat{g}(x)$ of the growth process $g(x)$. C. Transformation from measured distance $x$ (black curve, same as in panel A) to estimated time $t$ (time warping, magenta curve). Blue dashed lines are estimated peak positions. Magenta dashed lines are the warped peak positions. The horizontal axis is years (lower axis) and $m$ (distance on tusk; upper axis).
For each of the 12 pieces, we obtain a growth process $\hat{g}_{j}(x_{ij})$. The total number of observations is $n_{tot}=\sum_{j=1}^{12}n_{j}$. Re-enumerate the set of indices by concatenating chronologically
$ij\rightarrow k=i+\sum_{l=1}^{j}n_{l},\quad k=1,\ldots,n_{tot}.$ (24)
Assuming the signals can be glued together, without any gap in the timeline, we then construct an aggregated growth process:
$\hat{g}(x_{k})=\hat{g}(x_{ij})=\begin{cases}\hat{g}_{1}(x_{i1})&i\leq n_{1}\\ \hat{g}_{j-1}(x_{n_{j-1},j-1})+\hat{g}_{j}(x_{ij})&n_{j-1}<i\leq n_{j},\,j=2,\ldots,12\end{cases}.$ (25)
@@
$\textrm{Year}(k)=2010-\left(\hat{g}(x_{n_{tot}})-\hat{g}(x_{k})\right)/2\pi,\quad k=1,\ldots,n_{tot}.$ (26)
@@ 5.3 Age estimation
The quantity of interest is the age of the narwhal, which is equivalent to the number of cycles easily obtained from Eq. (5) and $\hat{g}(x_{n_{tot}})$. A confidence interval is obtained by bootstrapping (Algorithm 4). This is accomplished by performing parametric bootstrapping on each piece, and combine the bootstrapped signals randomly to produce $M=10^{7}$ realizations of aggregated growth processes $\hat{g}(x_{k})$, of which each has its own age estimate. The age estimate, including upper and lower confidence bounds, are summarized in table 2. For benchmarking, the table also lists the most common age estimator - manual counting of GLGs by our expert with proficiency in handling tusks - and the age estimate of a newer, modern approach using Carbon-14 measurements (Garde et al., 2024).
In the supplementary material, we present a Monte Carlo analysis to assess the coverage of our age estimates. While the nominal coverage is 95%, our results indicate a slight undercoverage. This discrepancy is likely due to the residual bootstrap occasionally producing multi-modal distributions of the age estimates. In such cases, the algorithm may converge to a solution that corresponds to a period doubling or halving.
Table 2: Age estimates with 95% confidence interval. * Based on radiocarbon dating using the bomb pulse (Zoppi et al., 2004) and Carbon-14 measurements from the narwhal tusk (Garde et al., 2024).
@@ 5.4 Model validation
We have overlain the fitted curves with the tusk pieces, to compare the identified peaks with the GLGs of the tusks. In Figure 9 we show two such examples. The laser line is highlighted in green and red, when inside and outside the dentin area, respectively. The positions of the peaks are marked with vertical red lines, and we can discern how these peaks appear to align with consecutive white lines, which presumably represent one GLG. A more formal model validation was carried out by inspecting the residuals (see Section 3.7) in Figure 10. The left plot shows a residual plot (raw residuals against fitted values) for each of the 12 fitted signals and the right plot displays QQ-plots, comparing the quantiles of the raw residuals to those from the theoretical standard normal distribution. The residual plots generally conform well to the homoscedasticity assumption and show only a few outliers, likely caused by rapid secondary cycles unique to our case study signals, which are not incorporated in Eq. (1). These cycles are likely associated with sub-annual accessory layers (Zhao, Matthews and Watt, 2025). In the QQ-plots, most residuals are close to the identity line, indicating that the central tendencies are well-captured by the model. However, skewed behavior in the tails, such as in piece Xs, suggests that the model may not fully capture the behavior of extreme values. We refer to supplementary material, where we more closely explore this departure from normality for pieces XIIx, XIu, and Xs.
@@
Figure 10: Model validation plots for the case-study signals. Residual plots (left) and $Q$-plots (right) for all 12 signals of tusk 956.
Estimator uncertainties within each tusk piece is visualized in Figure 11 using violinplots for all parameters, see Section 3.7. Among the $12$ signals of the case study, several estimators display a sizeable amount of variability and bias. For example, the one-lag autocorrelation $\rho$ is highly variable and shows some unsystematic bias (i.e. bias in no particular direction) across different pieces. The infinitesimal variance $\omega^{2}$ is also variable, but less prone to bias.
@@
Figure 11: Estimator variability of tusk 956. The $x$-axis represents the different tusk pieces (12 in total). For an estimate $\hat{\theta}$ of some model parameter $\theta$ we compute 100 bootstrap estimates $\tilde{\theta}$ and plot the relative differences $\Delta\theta:=(\tilde{\theta}-\hat{\theta})/\hat{\theta}$. The difference in phase $\Delta b$ is computed as the shortest distance on the unit circle, not in absolute difference, thus all differences are non-negative.
6 Discussion and outlook
The main attraction of the model is arguably the identification of the growth process $g(x)=\int_{s=0}^{x}\xi_{s}\,ds$, which effectively translates into identification of the timeline of the underlying hidden process (4) driving the changing periodicity. While the square-root diffusion process $(\xi_{x})_{x\geq 0}$ is slightly difficult to infer, likely caused by the integration of the process, the integrated process $g(x)$ appears to be well identified.
@@ 6 Discussion and outlook
It is important to consider the interplay between frequency and amplitude. For any time-varying quasi-periodic signal, we need to determine whether changes are primarily driven by variations in amplitude or frequency. In this study, we assumed that instantaneous jumps in the signal reflect stochastic shifts in frequency, while amplitudes vary slowly. This lets us first estimate and fix the amplitude envelope, then normalize to constant amplitudes before fitting our model. For the narwhal tusks, these assumptions imply that rapid changes in the signal occur due to changes in growth rate (on the order of days), whereas fluctuations in concentrations of elements in the environment are slow (on the order of months or years).
@@
While the presented model is both flexible and theoretically sound, the practical implementation still requires some choices which need a theoretical foundation, and should be tailored to the specific use. Some examples include selection of smoothing parameters in the initialization phase, runtime parameters (number of iterations inside Algorithm 1, particles in SMC and the threshold for the stopping criteria). These choices do not only assist the method in proper convergence, but are also important for controlling algorithmic runtime. These choices will serve the algorithm in producing more reliable estimates and guide the optimization procedure. We recommend to conduct preliminary simulation studies to guide the user for specific applications.
@@
[Acknowledgments] We thank the Greenlandic hunters for providing the narwhal tusk, Jeppe Møhl for sectioning the tusk, and Tonny Bernt Thomsen and Benjamin Heredia at the Geological Survey of Denmark and Greenland for performing the elemental analyses. The tusk was imported to Denmark from Greenland on CITES import permit no. IM 0822-296/18. Finally, we would like to thank Adam Gorm Hoffman, who was working with modelling and data analysis of a narwhal tusk in the early stage of the project.